### v1

@@ 3.4 Tree-width of $H$
In this section, we shall show that the output of Algorithm 1 is a triangle-free universally signable graph and conclude using Proposition 4. We begin by observing that $H$ is triangle-free.
@@
Assume for contradiction that, $H$ contains a triangle $T$. Then there must exist a cluster $S$ w.r.t $r$ and an edge $uv\in E(T)$ such that $\{u,v\}\subseteq S$. This contradicts 5(b). ∎
@@
Suppose $u,v$ lies in different clusters. For each $i\in\{1,2,3\}$, there exist an edge $x_{i}y_{i}\in E(P_{i})$ such that $x_{i}\in S$ and $y_{i}\in P\left(r,S\right)$. If $y_{1},y_{2},y_{3}$ are distinct vertices, we have a contradiction due to 3(a). Therefore, we can assume without loss of generality that $y_{1}=y_{2}=u$. This implies $x_{1}$ and $x_{2}$ are distinct vertices, and that $x_{1}\neq v,x_{2}\neq v$. However, now $S$ contains three vertices from an induced cycle in $H$ (formed by $V(P_{1})\cup V(P_{2})$), which contradicts Lemma 16.
Hence $u,v$ lies in the same cluster and let $U$ be the cluster that contains $P\left(r,S\right)$. Suppose there exists two vertices $\{v_{1},v_{2}\}\subseteq U$ such that $\{vv_{1},vv_{2}\}\subseteq E(T)$. Without loss of generality, assume that $v_{1}\in V(P_{1})$ and $v_{2}\in V(P_{2})$. Let $C$ be the induced cycle in $H$ induced by $V(P_{1})\cup V(P_{2})$. Then, for each $i\in\{1,2\}$ there exist edges $x_{i}y_{i}\in E(P_{i})$ such that $\{x_{1},x_{2}\}\subseteq U$ and $\{y_{1},y_{2}\}\subseteq S$. Applying Lemma 16 on $C$, $S$, and $U$, we have that $\{x_{1},x_{2}\}=\{v_{1},v_{2}\}$ and $\{y_{1},y_{2}\}=\{u,v\}$. However, this contradicts 4. Using similar arguments, we can show that there does not exist a cluster $S^{\prime}$ in $G$ such that $P\left(r,S^{\prime}\right)\subseteq S$ and $S^{\prime}$ contains two vertices that are adjacent to $v$ in $T$. Hence, it must be the case that there exists a vertex $v_{1}\in S\setminus\{u,v\}$ such that $v,v_{1}\in V(T)$. Hence, $S$ contains three vertices $u,v,v_{1}$ from an induced cycle in $H$, which contradicts Lemma 16. ∎
@@
Assume that $H$ contains a wheel whose rim is $C$ and the center is $u$. Let $S$ be the cluster that contains $u$ and let $u_{1},u_{2},u_{3}$ be three neighbors of $u$ in $C$. Suppose there exists a vertex $z\in\{u_{1},u_{2},u_{3}\}$ such that $d_{G}\left(r,z\right)=d_{G}\left(r,u\right)+1$ and $T$ be the cluster that contains $z$. Note that $P\left(r,T\right)\subseteq S$. Moreover, there must be a $z^{\prime}\in\{u_{1},u_{2},u_{3}\}\setminus\{z\}$ such that $z^{\prime}\notin T$. (Otherwise, $T$ will contain three vertices of $C$, which contradicts Lemma 16.) Suppose $d_{G}\left(r,z^{\prime}\right)=d_{G}\left(r,u\right)-1$. Then there exists two distinct $(z,z^{\prime})$-paths $Q_{1}\subseteq C,Q_{2}\subseteq C$, such that $V(Q_{1})\cap V(Q_{2})=\{z,z^{\prime}\}$. Observe that for each $i\in\{1,2\}$, there must be an edge $x_{i}y_{i}\in E(Q_{i})$ such that $x_{i}\in S$ and $y_{i}\in T$. Now, let $F\coloneq\{w\in P\left(r,T\right)\colon\exists v\in T,wv\in E(H)\}$. Observe that $\{x_{1},x_{2},u\}\subseteq F$, and hence $|F|\geq 3$, which contradicts 3(a). This implies $\{u_{1},u_{2},u_{3}\}\setminus\{z\}\subseteq S$, but this contradicts 5(a). ∎
Propositions 4, 17, 18 and 19 implies the following.
@@ 3.5 Completion of the proof
[…]
@@ 4 Conclusion
In this paper, combining the layering partition technique with the notion of strong isometric path complexity, we proved that $K_{2,3}$-induced minor-free graphs admit quasi-isometry with constant additive distortion to graphs with tree-width at most two. Minimizing the upper bound on the additive distortion is interesting. Since $K_{2,r}$-induced minor-free graphs and more generally $K_{2,r}$-asymptotic minor-free graphs [19] have bounded strong isometric path complexity, it would be interesting to study if these graph classes admit quasi-isometry with additive distortion to graphs with tree-width at most $r$. It would also be interesting to study if hereditary graph classes like even-hole free graphs [24], pyramid-only graphs, etc. admit quasi-isometry with additive distortion to graphs with tree-width at most two. A result of Albrechtsen et al. [2] implies that even-hole free and pyramid-only graphs admit quasi-isometry to graphs with tree-width at most two. However, both even-hole free graphs and pyramid-only graphs have unbounded strong isometric path complexity. Therefore, our technique cannot be used directly.

### v2

@@ 3.4 Tree-width of $H$
In this section, we shall show that the output of Algorithm˜1 is a triangle-free universally signable graph and conclude using Proposition˜5. We begin by observing that $H$ is triangle-free.
@@
Assume for contradiction that, $H$ contains a triangle $T$. Then there must exist a cluster $S$ w.r.t $r$ and an edge $uv\in E(T)$ such that $\{u,v\}\subseteq S$. This contradicts ˜5(b). ∎
@@
Suppose $u,v$ lies in different clusters. For each $i\in\{1,2,3\}$, there exist an edge $x_{i}y_{i}\in E(P_{i})$ such that $x_{i}\in S$ and $y_{i}\in P\left(r,S\right)$. If $y_{1},y_{2},y_{3}$ are distinct vertices, we have a contradiction due to ˜3(a). Therefore, we can assume without loss of generality that $y_{1}=y_{2}=u$. This implies $x_{1}$ and $x_{2}$ are distinct vertices, and that $x_{1}\neq v,x_{2}\neq v$. However, now $S$ contains three vertices from an induced cycle in $H$ (formed by $V(P_{1})\cup V(P_{2})$), which contradicts Lemma˜16.
Hence $u,v$ lies in the same cluster and let $U$ be the cluster that contains $P\left(r,S\right)$. Suppose there exists two vertices $\{v_{1},v_{2}\}\subseteq U$ such that $\{vv_{1},vv_{2}\}\subseteq E(T)$. Without loss of generality, assume that $v_{1}\in V(P_{1})$ and $v_{2}\in V(P_{2})$. Let $C$ be the induced cycle in $H$ induced by $V(P_{1})\cup V(P_{2})$. Then, for each $i\in\{1,2\}$ there exist edges $x_{i}y_{i}\in E(P_{i})$ such that $\{x_{1},x_{2}\}\subseteq U$ and $\{y_{1},y_{2}\}\subseteq S$. Applying Lemma˜16 on $C$, $S$, and $U$, we have that $\{x_{1},x_{2}\}=\{v_{1},v_{2}\}$ and $\{y_{1},y_{2}\}=\{u,v\}$. However, this contradicts ˜4. Using similar arguments, we can show that there does not exist a cluster $S^{\prime}$ in $G$ such that $P\left(r,S^{\prime}\right)\subseteq S$ and $S^{\prime}$ contains two vertices that are adjacent to $v$ in $T$. Hence, it must be the case that there exists a vertex $v_{1}\in S\setminus\{u,v\}$ such that $v,v_{1}\in V(T)$. Hence, $S$ contains three vertices $u,v,v_{1}$ from an induced cycle in $H$, which contradicts Lemma˜16. ∎
@@
Assume that $H$ contains a wheel whose rim is $C$ and the center is $u$. Let $S$ be the cluster that contains $u$ and let $u_{1},u_{2},u_{3}$ be three neighbors of $u$ in $C$. Suppose there exists a vertex $z\in\{u_{1},u_{2},u_{3}\}$ such that $d_{G}\left(r,z\right)=d_{G}\left(r,u\right)+1$ and $T$ be the cluster that contains $z$. Note that $P\left(r,T\right)\subseteq S$. Moreover, there must be a $z^{\prime}\in\{u_{1},u_{2},u_{3}\}\setminus\{z\}$ such that $z^{\prime}\notin T$. (Otherwise, $T$ will contain three vertices of $C$, which contradicts Lemma˜16.) Suppose $d_{G}\left(r,z^{\prime}\right)=d_{G}\left(r,u\right)-1$. Then there exists two distinct $(z,z^{\prime})$-paths $Q_{1}\subseteq C,Q_{2}\subseteq C$, such that $V(Q_{1})\cap V(Q_{2})=\{z,z^{\prime}\}$. Observe that for each $i\in\{1,2\}$, there must be an edge $x_{i}y_{i}\in E(Q_{i})$ such that $x_{i}\in S$ and $y_{i}\in T$. Now, let $F\coloneq\{w\in P\left(r,T\right)\colon\exists v\in T,wv\in E(H)\}$. Observe that $\{x_{1},x_{2},u\}\subseteq F$, and hence $|F|\geq 3$, which contradicts ˜3(a). This implies $\{u_{1},u_{2},u_{3}\}\setminus\{z\}\subseteq S$, but this contradicts ˜5(a). ∎
Propositions˜5, 17, 18 and 19 implies the following.
@@ 3.5 Completion of the proof
Recall that $G$ is a $K_{2,3}$-induced minor free graph, $r$ is an arbitrary vertex of $G$ (fixed in the beginning of 3), $H$ is the graph returned by Algorithm˜1. Let $c\coloneq 2470$ and $u,v$ be any two vertices of $G$. Due to Proposition˜7, the strong isometric path complexity of $G$ is at most $71$. Due to Lemmas˜7, 12 and 10, $d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\leq c$. Due to Lemmas˜17, 18 and 19 we know that $H$ is a universally signable graph, and therefore $K_{2,3}$-induced minor-free. Hence, the strong isometric path complexity of $H$ is at most $71$. Due to Lemmas˜7, 11 and 13, we have that $d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\leq c$. Hence, $\left|d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\right|\leq c$. Finally, Lemma˜20 asserts that $H$ has tree-width at most two. This completes the proof.
3.6 Remark on Corollary˜1
Any minimal cutset of universally signable graphs is either a clique or a stable set of size two (Proposition˜3). Therefore, when inputs of Algorithm˜1 are restricted to universally signable graphs the if condition of Algorithm˜1 can be modified as follows:
If $G\left[P\left(r,S\right)\right]$ contains at least three vertices or is an edge
This check can be done in $O(1)$-time, and the running time of Algorithm˜1 will be $O(n+m)$ where $n$ and $m$ are the number of vertices and edges of the input universally signable graph. Since the diameter of graphs of tree-width two can be found optimally in truly sub-quadratic time [7], we have the proof of Corollary˜1.
@@ 4 Conclusion
In this paper, combining the layering partition technique with the notion of strong isometric path complexity, we proved that $K_{2,3}$-induced minor-free graphs admit quasi-isometry with constant additive distortion to graphs with tree-width at most two. Minimizing the upper bound on the additive distortion is interesting. By tailoring the proof technique of Chakraborty et al. (Theorem 3, [8]) slightly and using Proposition˜1, the bound on the strong isometric path complexity of $K_{2,3}$-induced minor free can be improved. This would lead to an improvement on the bound on the additive distortion (to below $1000$). However, we believe the actual value of the optimum additive distortion consists of single digits. Providing a linear time algorithm to map $K_{2,3}$-induced minor-free graphs with additive distortion to graphs with tree-width at most two would be interesting.
Since $K_{2,r}$-induced minor-free graphs and more generally $K_{2,r}$-asymptotic minor-free graphs [22] have bounded strong isometric path complexity, it would be interesting to study if these graph classes admit quasi-isometry with additive distortion to graphs with tree-width at most $r$. It would also be interesting to study if hereditary graph classes like even-hole free graphs [29], pyramid-only graphs, etc. admit quasi-isometry with additive distortion to graphs with tree-width at most two. A result of Albrechtsen et al. [2] implies that even-hole free and pyramid-only graphs admit quasi-isometry to graphs with tree-width at most two. However, both even-hole free graphs and pyramid-only graphs have unbounded strong isometric path complexity. Therefore, our technique cannot be used directly.